OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). y = 2/3x − 10 y =2/3 x + 10 y = −3/2x − 10 y = −3/2x + 10

OpenStudy (p0sitr0n):

y=2/3x-1 perpendicular : -1/slope y=-3/2x+b 2=-3/2*-8+b b=-10 y=-3/2-10

OpenStudy (anonymous):

Thanks :D